OpenStudy (anonymous):

4x-y=24 and 2x+y=36. what values of x and y would make this true?

OpenStudy (freckles):

it is already setup for elimination add the equations together

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

they both have equal signs

OpenStudy (freckles):

\[\color{red}{4x-y}+\color{blue}{2x+y}=\color{red}{24}+\color{blue}{36}\]

OpenStudy (freckles):

if a=b and c=d then a+c=b+d

OpenStudy (anonymous):

oh